39. You bought a painting 10 years ago as an investment. You originally paid for $484,050, what was your annual return on investment? A) 47% B) 47% C) 1996 D) 12.8% Answer: $85,000 for it. If you sold it 40. What is the present value of S12.500 to be received 10 years from today? Assume a discount rate of 8% compounded annually and round to the nearest $10. A) $5,790 B) $11,574 C) $9,210 D) $17,010 Answer il. You wish to borrow $2,000 to be repaid in 12 monthly installments of $189.12. The annual interest rate A) 24% B) 8%. C)1896. D) 12%. Answer 42. annuities involve depositing money at the end of the period and allowing it to grow. A) Discount B) Compound C) Annuity due D) Both B and C Answer 43. Francis Peabody just won the $89,000,000 California State Lottery. The lottery offers the winner a choice of receiving the winnings in a lump sum or in 26 equal annual installments to be made at the beginning of each year. Assume that funds would be invested at 7.65%. Francis is trying to decide whether to take the lump sum or the annual installments. What is the amount of the lump sum that would be exactly equal to the present value of the annual installments? Round off to the nearest $1. A) $89,000,000 B) $38,163,612 C) $13,092,576 D) $41,083,128 Answer
Answer #1

39. ANSWER = C) 19%

Originally paid amount = $85000

Sold for = $484050

N = 10 years

PV = FV(1 + r)-n

85000 = 484050 * (1 + r) -10

85000 / 484050 = (1+ r)-10

0.175602 = (1+ r)-10

(0.175602)1/-10 = (1+ r)-10/-10

(0.175602)-0.10 = (1+ r)

1.1900 = 1 + r

1.1900 - 1 = r

0.1900 = r or r = 19.00%

40. Answer = a) $5,790

PV = FV * (1+r)-n

= $12500 * ( 1+ 0.08) -10

= $ 12500 * 0.463193

= $ 5790

41. ANSWER = A)24%

Borrowing amount = $ 2000

Monthly instalment = $ 189.12

n = 1 * 12 = 12

PVIFA = $ 2000 / 189.12

= 10.58

Refer to PVIFA Table, we get PVIFA(12) = 10.58 at 2%

Annual interest rate = 2% * 12 = 24 %

or PVIFA = [ 1 - (1+r)-n] / i

= [ 1 - (1 + 0.02)-12] / 0.02 = 10.58

42.

Answer = B)Compound: Annunities involve depositing money at the end of the period and allowing it to grow.